Use Cramer's rule to solve. If inconsistent or dependent, so state.
11x + 3y = 1
5x + 2y = 4

Answer by stanbon(75887)

11x + 3y = 1
5x + 2y = 4
---------
Coefficient Determinant = 22-15 = 7
----
x = (2-12)/7 = -10/7
y = (44-5)/7 = 39/7
========
Cheers.
Stan H.